OpenStudy (anonymous):

An amusement park ride consits of a large verticle cylinder that spins about its axis fast enough that a person inside is stuck to the wall and does not slide down when the floor drops away. The acceleration of gravity is 9.8m/s^2. Given g=9.8m/s^2, the coefficient of static friction between the wall and himan is 0.612, and the radius of the cylinder R=6.6m. For simplicity, neglect the person's depth and assume he or she is just a physical point on the wall. The person's speed is v=2piR/T. Where T is the rotation period of the cylinder (the time to complete a full circle). Find the maximum ro

OpenStudy (anonymous):

maximum ro...? Though did you attempt to draw a free body diagram?

OpenStudy (anonymous):

It actually is partiall drawn for me but i just labeled the forces onto the diagram

OpenStudy (anonymous):

find the maximum rotation i'm sorry.

OpenStudy (anonymous):

\[2\pi \sqrt{(coefficient of static friction)(radius) divided by 9.8m/s^2

OpenStudy (anonymous):

Well, for it to be just able to stick on the wall, that means there is no reaction force from the wall.  Solve the forces and find the speed on it. i.e \(F_c = \frac{v^2}{r}\)

OpenStudy (anonymous):

The velocity will slowly decrease because of friction.

OpenStudy (anonymous):

 this is the equation i first used but i got the wrong answer